# SBP-YOLO:A Lightweight Real-Time Model for Detecting Speed Bumps and Potholes

### v1

@@ Abstract
With the increasing demand for ride comfort from occupants of new energy vehicles, accurate and real-time detection of speed bumps and potholes has become critical for predictive suspension control. This paper presents SBP-YOLO, a lightweight detection framework built upon YOLOv11, optimized for embedded deployment. The proposed model integrates GhostConv for efficient computation, VoVGSCSPC for enhanced multi-scale feature representation, and a Lightweight and efficiency detection head (LEDH) to reduce the cost of early-stage feature processing. A hybrid training strategy combining NWD loss, knowledge distillation, and Albumentations-based weather augmentation is employed to improve detection accuracy and robustness, particularly for small and distant targets. Experimental results demonstrate that SBP-YOLO attains an 87.0% mAP, surpassing the baseline YOLOv11n by 5.8%, while achieving 139.5 FPS on the Jetson AGX Xavier platform following TensorRT FP16 quantization. This performance reflects substantial improvements in inference speed and computational efficiency. The findings validate the proposed method’s effectiveness for real-time road condition perception in intelligent suspension systems.
@@ 1 Introduction
With the rapid emergence of new energy vehicles (NEVs) as a mainstream transportation mode, consumers are placing increasingly higher demands on ride comfort and handling stability [1]. Compared to traditional internal combustion engine vehicles, electric vehicles (EVs) feature faster response times and smoother torque output, which pose greater challenges for the dynamic performance of suspension systems [2]. To ensure smooth driving and a comfortable experience under complex road conditions, real-time road condition perception has become critically important. Road surface irregularities such as speed bumps and potholes are primary factors affecting vehicle dynamics and ride comfort [3], as these disturbances are transmitted through the suspension system to the vehicle body, directly impacting passenger experience. The key to advanced suspension control lies in the anticipatory perception and accurate prediction of road disturbances, including the classification of stochastic road profiles and the identification of impulse-type irregularities [4]. Intelligent suspension systems, leveraging previewed road information, can dynamically adjust stiffness and damping to adapt to complex road environments. This enables effective suppression of vehicle body vibrations and attitude changes, thereby enhancing driving stability and safety [5, 6]. In EVs in particular, the rapid torque response of electric motors imposes greater demands on the perception and reaction speed of suspension systems. This is particularly critical when encountering sudden road irregularities such as potholes and speed bumps, which require swift and precise adaptation. Consequently, more tightly coordinated integration with the overall vehicle dynamics is essential to maintain both ride comfort and driving safety.
Real-time detection of road surface anomalies remains a technical challenge. Traditional methods rely on vehicle dynamics, image processing, or costly sensor setups [7, 8]. Dynamics-based approaches analyze acceleration and vibration signals under varying road conditions to identify anomalies such as speed bumps and potholes. For instance, Celaya-Padilla et al. [9] introduced a sensor-fusion-based method for speed bump detection, employing gyroscope, accelerometer, and GPS data, with a logistic model optimized via cross-validation and genetic algorithms. Salman et al. [10] used smartphone sensors with CNN–LSTM models to detect and classify speed bumps with up to 98.92% accuracy. Similarly, Yin et al. [11] proposed a smartphone-based approach utilizing acceleration and GPS data, where signal preprocessing and an Extracted Data Feature Filter (EDFF) were applied to effectively differentiate between potholes and speed bumps. Aguilar-González et al. [12] developed a random forest classifier using tri-axial vibration data for road event detection. Despite their effectiveness, these methods are typically task-specific and computationally expensive, lacking the predictive capability required for real-time suspension control in intelligent vehicles.
In recent years, the rapid development of deep learning has significantly advanced object detection techniques based on computer vision, providing new solutions for identifying road surface anomalies such as speed bumps and potholes. Compared with conventional methods relying on accelerometer signals, deep learning-based vision approaches enable proactive perception of road surface conditions, allowing earlier and more accurate detection of anomalies for suspension predictive control. The YOLO (You Only Look Once) series is widely utilized in road surface anomaly detection [13], offering a favorable trade-off between detection accuracy and inference speed, which is critical for real-time perception in intelligent vehicles. For example, Park et al. [14] conducted a comparative study of three YOLO variants, namely YOLOv4, YOLOv4-tiny, and YOLOv5s, and concluded that YOLOv4-tiny provided the most favorable trade-off between real-time performance and detection accuracy for pothole recognition. ASAD et al. [15] implemented a lightweight Tiny-YOLOv4 model on a Raspberry Pi device for edge-based pothole detection, achieving a detection accuracy of 90% at 31.76 FPS. Khan et al. [16] introduced a YOLOv8-based real-time pothole detection method for autonomous vehicles, outperforming YOLOv5 on public datasets through data augmentation. To address challenges caused by variable lighting and adverse weather conditions, Bučko et al. [17] constructed a dataset containing road images captured under different environmental scenarios and validated the performance of the YOLOv3 algorithm accordingly. Furthermore, Bhavana et al. [18] proposed POT-YOLO based on YOLOv8, capable of detecting various pavement anomalies such as cracks, oil stains, patches, and cobblestones, achieving a detection accuracy of 97.6%. These studies highlight the effectiveness of YOLO-based detection algorithms in real-time identification of road anomalies under complex driving conditions.
@@
    LEDH is proposed to alleviate the overhead of the P2 detection layer, maintaining accuracy for small, distant targets while reducing head complexity from dense, high-dimensional convolutions.
@@
    GhostConv modules are applied to the backbone and neck for efficient feature extraction, while VoVGSCSPC improves robustness across varying road and environmental conditions.
@@
    To improve detection of small and distant targets in dynamic environments, a comprehensive training strategy is adopted, integrating NWD Loss [19] for accurate localization, BCKD [20] for knowledge distillation, and Albumentations-based [21] augmentation to simulate motion blur, lighting changes, and adverse weather, enhancing robustness and generalization.
  • $\bullet$
    Quantized to FP16 and optimized with TensorRT, the model achieves 139.5 FPS inference speed on the NVIDIA Jetson AGX Xavier, meeting strict latency demands for real-time road detection and enabling reliable embedded deployment.
@@ 2 Methods
2.1 SBP-YOLO
YOLOv11 [22] is the latest advancement in the YOLO series of object detection frameworks, integrating the C3k2 block, SPP-Fast module, and the parallel spatial attention C2PSA block, which collectively enhance feature extraction and detection performance. The series includes multiple variants ranging from lightweight to large-scale models; this study selects the lightweight YOLOv11n as the baseline to meet the real-time and computational constraints of embedded platforms. However, existing models still face challenges in handling multi-scale targets and complex backgrounds, especially in accurately detecting small and distant road anomalies such as potholes and speed bumps. To mitigate these limitations, this paper proposes the SBP-YOLO model Fig. 1: GhostConv modules replace standard convolutions (SC) in the deeper network layers to reduce computational cost while preserving strong feature extraction; the VoVGSCSPC module enhances multi-scale feature representation; a dedicated P2 detection layer and LEDH are designed to improve small-object detection; The overall network architecture is annotated with output feature map dimensions, such as (320, 320, 16), representing height, width, and channel depth, to clearly illustrate the scale of features at each stage.
@@ 2.1 SBP-YOLO
2.2 GhostConv
Standard convolutional (SC) modules often generate redundant feature maps, resulting in high computational and memory costs. To mitigate this, the GhostConv module from GhostNet [23] generates additional “ghost” feature maps through inexpensive operations, effectively reducing complexity while preserving representational capacity. As illustrated in Fig. 2, GhostConv employs a two-step process: first, intrinsic feature maps are produced by SC with small kernels; subsequently, $\times 5$ 5 depthwise convolutions (DWConv) expand these to approximate the full feature set. The outputs from both steps are fused, delivering comparable results to SC but with substantially fewer parameters and reduced computation. By applying DWConv to only half of the channels, GhostConv efficiently enlarges the receptive field and lowers computational overhead.
Figure 2: GhostConv structure diagram
@@ 2.2 GhostConv
Let $c_{1}$ and $c_{2}$ be the input and output channels, $k$ the kernel size, and $H\times W$ the input feature map dimensions. Ignoring bias, the parameter count $P_{\text{Conv}}$ and FLOPs $F_{\text{Conv}}$ of a SC are:
$P_{Conv}=c_{1}c_{2}k^{2},\quad F_{Conv}=2c_{1}c_{2}HWk^{2}.$ (1)
In GhostConv, the main convolution produces half of the output channels, $\frac{c_{2}}{2}$, followed by a cheap convolution generating the remainder. The total parameter count $P_{g}$ and computational cost $F_{g}$ are:
$P_{g}=c_{1}\frac{c_{2}}{2}k_{m}^{2}+\frac{c_{2}}{2}k_{c}^{2},\quad F_{g}=c_{1}c_{2}HWk_{m}^{2}+c_{2}HWk_{c}^{2},$ (2)
where $k_{m}$ and $k_{c}$ denote the kernel sizes of the main and cheap convolutions, respectively. For example, at SBP-YOLO’s 7th layer, the GhostConv module reduces theoretical computation to 51.1% and parameter count by 48.9% compared to SC, markedly decreasing model complexity and size.
2.3 VoVGSCSPC
Depthwise separable convolution (DSConv) reduces computational cost and parameters by decomposing SC into depthwise convolution (DWConv) for spatial filtering and pointwise convolution (PWConv) for channel mixing. However, since spatial and channel information are processed separately in DWConv, inter-channel information exchange is limited, leading to feature isolation within groups. To mitigate this issue, ShuffleNets [24] introduced the channel shuffle operation to enable cross-channel interaction, while MobileNet [25] employed extensive $\times 1$ 1 dense convolutions to fuse channel information. Nevertheless, the dense $\times 1$ 1 convolutions increase the computational burden. Although channel shuffle and other lightweight operations improve efficiency, their performance remains inferior to SC.
GSConv [26] addresses these limitations by integrating the advantages of GhostConv with a hardware-friendly, efficient channel shuffle mechanism. This design overcomes the information isolation problem inherent in DWConv, enabling efficient intra-group feature flow and enhancing feature representation and overall model performance. GSConv not only reduces computational complexity but also improves cross-group feature fusion, achieving accuracy comparable to or exceeding SC in lightweight networks. This makes GSConv a promising solution for high-performance visual tasks on mobile and embedded platforms.
@@ 2.3 VoVGSCSPC
VoVGSCSPC extends GSConv with a cross-stage design, splitting the input into two branches: one with a convolution followed by a GSBottleneck, and another with a DWConv residual branch. Their outputs are fused as
@@
$VoVGSCS{P}_{\mathrm{out}}=\alpha\bigl{(}\mathrm{Concat}(GSB_{\mathrm{out}},\alpha(X_{C_{1}}))\bigr{)}$ (5)
where $GSB_{\mathrm{out}}$ and $VoVGSCS{P}_{\mathrm{out}}$ denote intermediate and final outputs, respectively.
@@ 2.4 Lightweight and efficiency detection head (LEDH)
In suspension preview control, road anomalies such as speed bumps and potholes often appear as small, low-saliency objects at a distance, posing typical small object detection challenges. Although the inclusion of a shallow P2 detection layer in YOLOv11 enhances detection performance for such targets, it also introduces considerable computational overhead, hindering deployment on resource-constrained edge devices. In response to this challenge, we propose the LEDH, which preserves accuracy on small targets while significantly reducing computational complexity, enabling real-time inference in embedded systems.
Figure 4: Lightweight and efficiency detection head (LEDH) structure Inputs P2, P3, P4, and P5 go through 3 × 3 GroupConv and two Conv2D layers, then concat into two branches for classification and regression to yield predictions
As shown in Fig. 4, LEDH adopts a decoupled architecture with parallel branches for classification and regression. To enhance feature expressiveness with minimal cost, each branch includes two stacked $\times 3$ 3 grouped convolutions with group count $g=c_{i}/16$, applied to the input feature $\mathbf{x}_{i}$ at scale $i$:
$F_{i}$ $=\mathrm{Conv}_{g}\big{(}\mathrm{Conv}_{g}(\mathbf{x}_{i})\big{)},$ (6)
$H_{i}$ $=\mathrm{Concat}\left[\mathrm{Conv}_{1\times 1}^{4r}(F_{i}),\ \mathrm{Conv}_{1\times 1}^{n_{c}}(F_{i})\right]$
where $F_{i}$ denotes the enhanced feature, while $H_{i}$ represents the final prediction output, consisting of box regression logits and class probabilities. The terms $n_{c}$ and $r$ refer to the number of classes and the DFL bin count, respectively. LEDH processes multi-scale features (P2–P5) from the backbone, enriching both spatial resolution and semantic depth through grouped convolutions. This structure reduces parameter count and FLOPs while preserving receptive field and non-linear capacity. Final predictions are produced via lightweight $\times 1$ 1 convolutions for classification and bounding box regression. The design balances detection accuracy and computational efficiency, particularly benefiting small-object detection tasks under edge deployment constraints.

### v2

@@ Abstract
Reliable and real-time detection of road speed bumps and potholes is crucial for anticipatory perception in advanced suspension systems, enabling timely and adaptive damping control. However, achieving high accuracy and efficiency on embedded platforms remains challenging due to limited computational resources and the small scale of distant targets. This paper presents SBP-YOLO, a lightweight and high-speed detection framework tailored for bump and pothole recognition. Based on YOLOv11n, the model integrates GhostConv and VoVGSCSPC modules into the backbone and neck to reduce computation while enhancing multi-scale semantic features. To improve small-object detection, a P2-level branch is introduced, with a Lightweight and efficient detection head (LEDH) mitigating the added computational overhead without compromising accuracy. A hybrid training strategy combining NWD loss, BCKD-based knowledge distillation, and Albumentations-driven augmentation further enhances localization precision and robustness. Experiments show that SBP-YOLO achieves 87.0% mAP, outperforming the YOLOv11n baseline by 5.8%. After TensorRT FP16 quantization, it runs at 139.5 FPS on Jetson AGX Xavier, delivering a 12.4% speedup over the P2-enhanced YOLOv11. These results validate the effectiveness of the proposed method for fast and low-latency road condition perception in embedded suspension control systems.
@@ 1 Introduction
Adjustable suspension systems, such as magnetorheological semi-active [1, 2] and fully active suspensions [3], offer real-time tunability of damping and stiffness to enhance ride comfort and handling stability [4]. However, their performance largely depends on timely and accurate perception of upcoming road conditions [5]. In the absence of reliable preview information, these systems are limited to reactive control, making it difficult to fully exploit their potential in suppressing body vibrations and improving vehicle dynamics. Among various road disturbances, speed bumps and potholes represent common yet critical anomalies that directly impact ride quality [6]. Early detection of such features enables predictive suspension control, allowing damping strategies to be adjusted in advance[7, 8, 9]. Existing methods for road anomaly detection rely on vehicle dynamics, vibration analysis, or multi-sensor fusion (e.g., accelerometers, gyroscopes, and GPS) [10, 11, 12]. While effective, these approaches are often task-specific, computationally intensive, or limited by high hardware costs, making them unsuitable for real-time embedded deployment in advanced suspension systems.
Camera-based visual perception offers a cost-effective and information-rich alternative to traditional sensor-based approaches for road condition monitoring. Deep learning has substantially advanced vision-based anomaly detection by enabling proactive identification of road defects such as speed bumps and potholes. Compared to acceleration signal analysis, vision-based methods provide earlier and more reliable detection, offering critical input for real-time suspension control. The YOLO (You Only Look Once) series has further improved the trade-off between detection accuracy and inference speed [13]. Bučko et al. [14] evaluated YOLOv3 for pothole detection under diverse lighting and weather conditions, and compared it with Sparse R-CNN. Their results showed that while Sparse R-CNN offered slightly higher robustness in extreme visual degradation, YOLOv3 remained competitive due to its lower hardware requirements and stable performance across most conditions. Subsequent versions introduced structural improvements: YOLOv4 adopted CSPDarknet and PANet for enhanced feature aggregation, while its lightweight variant YOLOv4-tiny was deployed by Park et al. [15] for pothole detection with superior real-time capability. ASAD et al. [16] implemented a lightweight YOLOv4-tiny model on a Raspberry Pi device for edge-based pothole detection, achieving a detection accuracy of 90% at 31.76 FPS. YOLOv5 [17] incorporated anchor-free heads and auto-learned anchor boxes, whereas YOLOv8 [18] leveraged decoupled heads and improved data augmentation techniques. Khan et al. [19] further optimized YOLOv8 with customized augmentations for autonomous driving scenarios. Building on these, Bhavana et al. [20] proposed POT-YOLO based on YOLOv8, achieving 97.6% accuracy in detecting complex pavement defects. More recently, YOLOv11 [21] introduced the C3k2 residual block and C2PSA attention mechanism to enhance multiscale feature representation. Its modular architecture spans from lightweight to large-scale variants, supporting flexible deployment across both embedded and high-performance platforms.
@@
    A lightweight and efficiency detection head (LEDH) is introduced to reduce the computational overhead of the P2 layer, maintaining accuracy for small and distant targets. Combined with GhostConv modules and VoVGSCSPC blocks in the backbone and neck, the proposed architecture enhances feature extraction efficiency and robustness under diverse road and environmental conditions.
@@
    A comprehensive training strategy is adopted to improve detection performance in dynamic scenarios, incorporating NWD Loss [22] for precise localization, BCKD [23] for knowledge distillation, and Albumentations-based [24] data augmentation to simulate motion blur, lighting variations, and adverse weather conditions.
@@
    The model is quantized to FP16 and optimized using TensorRT, achieving an inference speed of 139.5 FPS on the NVIDIA Jetson AGX Xavier, thereby meeting real-time requirements for embedded road condition detection.
The remainder of this paper is organized as follows: Sect. 2 presents the proposed model architecture and the improved loss functions. Sect. 3 details the experimental setup and evaluation results. Sect. 4 discusses the model’s inference performance on the embedded platform. Finally, conclusions are drawn in Sect. 5.
@@ 2 Methods
2.1 SBP-YOLO Architecture
@@
This study focuses on the requirements of intelligent suspension systems, where accurate detection of road surface anomalies such as potholes and speed bumps is essential for predictive control. To address the constraints of limited computational resources and real-time inference, we adopt YOLOv11n as the baseline and introduce a series of structural improvements to form the proposed SBP YOLO architecture, as illustrated in Fig. 1.
The backbone network is enhanced by replacing conventional convolutional layers with GhostConv modules in deeper stages. This substitution significantly reduces computational cost while maintaining sufficient feature extraction capability. To improve the representation of features across different spatial resolutions, a new module named VoVGSCSPC is introduced. This design integrates densely connected patterns inspired by VoVNet with a GSConv-based bottleneck structure, and it is applied in both the backbone and the detection head to enhance semantic feature propagation.
To increase sensitivity to small and distant objects, an additional early-stage P2 detection layer is incorporated. This layer allows the model to capture fine spatial details by leveraging features from the shallow stages of the network. In contrast to the standard three-level detection structure (P3 to P5), SBP YOLO employs a four-level detection strategy that includes feature maps from P2 (160 $\times$ 160), P3 (80 $\times$ 80), P4 (40 $\times$ 40), and P5 (20 $\times$ 20). The detection head also integrates a lightweight and efficient design, referred to as LEDH, which further reduces computational overhead while preserving detection accuracy. All feature maps are annotated with their spatial dimensions, such as (320, 320, 16), where the three values represent the height, width, and number of channels, respectively. This notation provides a clear overview of the scale and structure of each layer.
Overall, SBP YOLO extends the baseline YOLOv11n by incorporating GhostConv modules, the VoVGSCSPC structure, a dedicated early detection branch, and the LEDH. These enhancements are designed to improve the model’s ability to capture fine spatial details, enhance feature representation across multiple scales, and reduce computational complexity. Such architectural improvements aim to enable accurate and efficient detection of road surface anomalies, supporting deployment in intelligent vehicle control and predictive suspension systems.
2.2 GhostConv Module
In practical road anomaly detection, identifying small and sparsely distributed targets—such as distant speed bumps and surface depressions—remains challenging due to their limited spatial coverage and low contrast with the surrounding environment. Effective detection of such anomalies requires a large receptive field while maintaining computational efficiency, which is particularly important for embedded platforms with constrained processing resources.
To address this issue, we incorporate the Ghost convolution (GhostConv) module from GhostNet [25] as a lightweight substitute for standard convolution (SC). GhostConv follows a two-stage strategy: an initial small-kernel convolution extracts primary features, which are then enriched by a depthwise convolution using a larger kernel (e.g., $5\times 5$) to generate supplementary "ghost" features. The depthwise convolution operates on half of the intermediate channels, allowing expansion of the receptive field with minimal computational overhead.
Integrating GhostConv into the backbone of SBP-YOLO enhances spatial feature representation while maintaining low complexity. This enables more accurate recognition of fine-grained road anomalies and supports efficient inference on low-power devices. As shown in Fig. 2, GhostConv significantly reduces the number of parameters and floating point operations (FLOPs) without sacrificing feature representation capability.
@@ 2.2 GhostConv Module
Let $c_{1}$ and $c_{2}$ represent the input and output channels, $k$ the kernel size, and $H\times W$ the spatial resolution. The parameter count and FLOPs for standard convolution are given by:
$P_{\text{SC}}=c_{1}c_{2}k^{2},\quad F_{\text{SC}}=2c_{1}c_{2}HWk^{2}.$ (1)
For GhostConv, the initial convolution uses kernel size $k_{m}$, and the depthwise convolution uses $k_{c}$. The corresponding parameter count and FLOPs are:
$P_{g}$ $=c_{1}\cdot\frac{c_{2}}{2}\cdot k_{m}^{2}+\frac{c_{2}}{2}\cdot k_{c}^{2},$ (2)
$F_{g}$ $=2c_{1}\cdot\frac{c_{2}}{2}\cdot HWk_{m}^{2}+2\cdot\frac{c_{2}}{2}\cdot HWk_{c}^{2}.$ (3)
The cost ratios relative to standard convolution are:
$\frac{P_{g}}{P_{\text{SC}}}=\frac{k_{m}^{2}+\frac{k_{c}^{2}}{c_{1}}}{2k^{2}},\quad\frac{F_{g}}{F_{\text{SC}}}=\frac{k_{m}^{2}+\frac{k_{c}^{2}}{c_{1}}}{k^{2}}.$ (4)
Under typical settings (e.g., $c_{1}\in[64,256]$, $k_{m}=3$, $k_{c}=5$), GhostConv reduces parameters and FLOPs to approximately 48%–55% of those of standard convolution. This demonstrates its effectiveness for efficient deployment in embedded detection frameworks. Moreover, Fig. 2(c) shows visualized feature maps from the 5th and 7th layers in the backbone. The GhostConv-enhanced model exhibits stronger responses to distant small-scale targets, highlighting its improved spatial awareness in complex road scenes.
Figure 2: Illustration of the Ghost module design and its impact on feature activation. (a) Ghost module; (b) GhostConv architecture; (c) Comparison of backbone activation heat maps at the 5th and 7th convolutional layers.
2.3 VoVGSCSPC Module
@@
Depthwise separable convolution (DSConv) reduces computational cost by decomposing standard convolution (SC) into a depthwise convolution (DWConv) for spatial filtering and a pointwise convolution (PWConv) for channel mixing. However, as spatial and channel operations are decoupled, DWConv suffers from limited inter-channel interaction, leading to feature isolation across groups. To address this, ShuffleNet [26] introduced channel shuffle to enable cross-channel information flow, while MobileNet [27] adopted extensive $1\times 1$ convolutions for dense channel fusion. Despite their effectiveness, these methods either suffer from reduced expressiveness or increased computational overhead, and often underperform SC in complex tasks.
GSConv [28] mitigates these drawbacks by combining the efficiency of GhostConv with a hardware-friendly channel shuffle mechanism. It enables effective intra- and inter-group feature communication, alleviating the isolation issue inherent in DWConv. By enhancing feature representation with minimal overhead, GSConv achieves accuracy comparable to or better than SC in lightweight models, making it well-suited for mobile and embedded visual applications.
@@
VoVGSCSPC [29] extends GSConv with a cross-stage architecture, splitting the input into two branches: one with a convolution followed by a GSBottleneck, and another with a DWConv residual branch. Their outputs are fused as
@@
$VoVGSCSP{C}_{\mathrm{out}}=\alpha\bigl{(}\mathrm{Concat}(GSB_{\mathrm{out}},\alpha(X_{C_{1}}))\bigr{)}$ (7)
where $GSB_{\mathrm{out}}$ and $VoVGSCSP{C}_{\mathrm{out}}$ denote intermediate and final outputs, respectively.
@@ 2.4 Lightweight and efficiency detection head (LEDH)
In suspension preview control, road anomalies such as speed bumps and potholes often appear as small, low-saliency objects at a distance, posing challenging small-object detection problems. Although YOLOv11 improves performance by introducing a shallow P2 detection layer, each detection head still employs two consecutive $3\times 3$ convolutions before the classification and regression branches, leading to considerable computational overhead and hindering deployment on embedded platforms.
To address this limitation, we propose the LEDH, a lightweight detection head that reduces redundancy while maintaining accuracy on small targets. As shown in Fig. 4, for each scale $i$, the input feature ${c}_{i}$ is first transformed by two stacked grouped convolutions (GroupConv) with group count $g=c_{i}/16$:
$F_{i}$ $=\mathrm{Conv}_{g}\big{(}\mathrm{Conv}_{g}({c}_{i})\big{)},$ (8)
$H_{i}$ $=\mathrm{Concat}\left[\mathrm{Conv}_{1\times 1}^{4r}(F_{i}),\ \mathrm{Conv}_{1\times 1}^{n_{c}}(F_{i})\right],$
where $F_{i}$ denotes the shared transformed feature map and $H_{i}$ the detection head output containing bounding box regression logits and class probabilities. Here, $n_{c}$ and $r$ are the number of classes and the Distribution Focal Loss (DFL) bin count, respectively.
Unlike YOLOv11, which applies separate convolutional blocks to the classification and regression branches, LEDH introduces a shared transformation stage with GroupConv, so that the enhanced feature representation is jointly utilized by both branches. This design eliminates redundant computation, significantly reducing parameters and FLOPs, while preserving spatial detail and semantic abstraction across scales. Subsequent lightweight $1\times 1$ convolutions produce the final predictions, yielding an efficient detection head well suited for real-time edge deployment.
Figure 4: Structure of the proposed LEDH. Multi-scale features (P2–P5) are transformed by two stacked $3\times 3$ GroupConv layers and then fed into parallel $1\times 1$ convolution branches for classification and regression.